# White-box Multimodal Jailbreaks Against Large Vision-Language Models

### v1

@@ 4.1.1 Datasets
4.1.2 Metrics
Here we describe the metrics used in our evaluations. Attack Success Rate (ASR). Given the tendency of keyword-based automated assessments to overinflate ASRs, we implement strict manual evaluation on both the Advbench’s (Zou et al., 2023) harmful behaviors and the VAJM (Qi et al., 2023) evaluation set. Attacks that generate relevant and useful harmful output are defined as successful, with others considered failures. Toxicity Rate. To measure the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we utilize the Perspective API and the Detoxify classifier (Hanu & Unitary team, 2020), both of which can calculate toxicity scores for a set of six toxic attributes. Scores range from 0 (least toxic) to 1 (most toxic). For each attribute, we calculate the proportion of texts whose toxicity scores exceed a 0.5 threshold.
4.1.3 Implementation Details
To evaluate our attack methodology, we employ the Vicuna-13B version of MiniGPT-4 (Zhu et al., 2023). It is built on the frozen Vicuna-13B (Chiang et al., 2023) backbone model, an aligned LLM derived from LLaMA (Touvron et al., 2023). For the experimental setup, we adopt the configuration utilized in VAJM (Qi et al., 2023), specifying a batch size $b$ of 8, a step size $\alpha$ of 1, and an iteration count $N_{1}$ of 5000. Furthermore, we set the image-text attack ratio $r$ to 10, the length of the adversarial text suffix to 20 tokens, the number of candidates $n$ to 250, and the iteration count $N_{2}$ to 2000 by default. We use a single A100 GPU with 80GB of memory in all experiments.
4.2 Comparison With Unimodal Attacks
@@
Table 1: Comparison of Train ASR (%) and Test ASR (%) on Advbench’s (Zou et al., 2023) harmful behaviors. GCG-V refers to the visual version of GCG.
@@
Table 2: Comparison of average ASR (%) measured on the VAJM (Qi et al., 2023) evaluation set across 4 categories of harmful instructions.
ASR(%) Identity Attack Disinfo Violence/Crime X-risk Without Attack 30.8 53.3 57.3 33.3 GCG (Zou et al., 2023) 49.2 48.9 57.3 40.0 VAJM (Qi et al., 2023) 81.5 82.2 85.3 60.0 GCG-V 66.2 64.4 84.0 6.7 UMK (Ours) 87.7 95.6 98.7 46.7
Table 3: Percentages (%) of outputs that display speciffc toxic attributes evaluated by the Perspective API and Detoxify classifier (Hanu & Unitary team, 2020). These outputs are generated on the ‘challenging’ subset from RealToxicityPrompts benchmark (Gehman et al., 2020). ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
@@
Table 4: Comparison of Train ASR (%) and Test ASR (%) for ablation studies on Advbench’s (Zou et al., 2023) harmful behaviors.
@@
To further broaden our evaluation, we utilize the challenging subset of RealToxicityPrompts benchmark (Gehman et al., 2020), which includes 1,225 text prompts designed to trigger toxic continuations. Given that this is a text continuation task, we exclude the GCG method (Zou et al., 2023), which focuses on generating text suffixes, from our comparison. Similarly, for our approach, we utilize only the adversarial image prefix. We pair each adversarial image with text prompts from the dataset as input, and then measure the toxicity of the outputs using both the Perspective API and the Detoxify classifier [31], each capable of calculating toxicity scores for a set of six toxic attributes. For each attribute, we calculate the ratio of the generated texts with scores exceed the threshold of 0.5. As shown in Table 3, VAJM (Qi et al., 2023) and GCG-V display different performances across two evaluation methods. Under the Perspective API, GCG-V achieves a 71.98% toxicity rate, higher than VAJM’s 67.17%. Conversely, with the Detoxify classifier (Hanu & Unitary team, 2020), VAJM records a 61.96% rate, exceeding GCG-V’s 56.34%. Although VAJM achieves results superior to those of our method in certain categories such as Identity Attack, this is attributable to its optimization objective being primarily focused on generating harmful statements against gender, race, and humanity. In contrast, our method’s optimization objectives are more comprehensive. Although our method only utilizes adversarial image prefix in this experiment, it achieves overall best results, recording Any* toxicity rates of 76.98% and 68.70% under the Perspective API and Detoxify classifier tests, respectively. These results significantly outperform the previous best rates of 71.98% from GCG-V and 61.96% from VAJM. This demonstrates the effectiveness of our dual optimization objective strategy: even though the second phase is aimed at guiding the model to generate affirmative responses, it successfully enhances the toxic semantics of the adversarial image. Moreover, the joint attack of the text-image multimodalities may have uncovered a more optimal solution space for the adversarial image.
4.3 Ablation Studies
@@
4.4 Understanding the Role of Text Attack in Multimodal Strategy
Figure 4: Comparative Analysis of Loss Before and After Text Attack. The X-axis represents ‘Steps’, while the Y-axis denotes ‘Loss’.
@@ 4.4 Understanding the Role of Text Attack in Multimodal Strategy
Figure 5: Overview of the image-image attack strategy. We adopt the same dual optimization objectives as used in the text-image attack.
Figure 6: Challenging examples of our proposed multimodal attack strategy that circumvents MiniGPT-4’s (Zhu et al., 2023) safety mechanisms, showcasing generated content that promotes harmful behaviors and makes contentious political comparisons.
In this subsection, we explore the role of text in the text-image multimodal attack strategy. Figure 4 illustrates the variance in batch loss before and after text attack throughout the execution of our attack methodology. For ease of analysis, we employ linear regression models to fit the curve. Observations at different steps indicate a significant reduction in batch loss after the text attack. As the number of steps increases, an overall downward trend in loss can be observed, demonstrating the efficacy of our method. Notably, even in the later stages of the attack, the reduction in batch loss before and after the text attack remains evident, emphasizing the pivotal contribution of text in amplifying the effectiveness of our strategy.
To better understand the role of text in our proposed text-image multimodal attack strategy, we introduce an image-image attack for comparative analysis. As illustrated in Figure 5, during the attack, we input two adversarial images and optimized them simultaneously. In the case of MiniGPT-4 (Zhu et al., 2023), the adversarial images we input each occupies 32 tokens, while the adversarial text suffix proposed in our attack utilizes only 20 tokens. Given that the textual space being discrete and denser compared to the visual space, adversarial attacks in the textual domain are generally more demanding. Intuitively, one might anticipate that image-image attacks, which are optimized over a larger volume of tokens, would yield superior results. However, this is not the case. In Table 5, we report the average ASR for both the image-image attack and the proposed text-image attack on the VAJM (Qi et al., 2023) evaluation set. It is evident that the efficacy of the image-image attack is significantly inferior to that of our attack. We observe that the image-image attacks are more likely to trigger the model to repeat affirmative responses twice or to cease output after the affirmative response. We believe this is because the optimization objectives for both adversarial images in the second phase are to maximize the probability of the model’s affirmative responses. In this context, the semantics of the two images tend to overshadow that of the text input. Consequently, during the testing phase, two adversarial images containing the same semantics are more likely to cause the model to duplicate affirmative reactions or only produce affirmative responses. On the other hand, this phenomenon can also be attributed to the training strategies employed by VLMs. Since VLMs are typically trained on image-text pairs, they tend to experience a performance degradation when encountering out-of-distribution inputs such as image-image-text trios.
4.5 Qualitative Analysis
@@
In Figure 6, we present two challenging attack examples. Our text-image multimodal attack strategy can effectively circumvent the safety alignments of MiniGPT-4 (Zhu et al., 2023), generating content such as messages that encourage self-harm or suicide and comments that comparing politicians to Nazis.
@@
Table 5: Comparative analysis of average ASR (%) between image-image and text-image attacks on the VAJM (Qi et al., 2023) evaluation set.
ASR(%) Identity Attack Disinfo Violence/Crime X-risk Image-Image Attack 78.5 68.9 85.3 6.7 Text-Image Attack 87.7 95.6 98.7 46.7
5 Conclusion

### v2

@@ 4.1.1. Datasets
4.1.2. Metrics
Here we describe the metrics used in our evaluations. Attack Success Rate (ASR). Given the tendency of keyword-based automated assessments to overinflate ASRs, we implement strict manual evaluation on both the Advbench’s (Zou et al., 2023) harmful behaviors and the VAJM (Qi et al., 2023) evaluation set. Attacks that generate relevant and useful harmful output are defined as successful, with others considered failures. Toxicity Rate. To measure the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we utilize the Perspective API and the Detoxify classifier (Hanu and Unitary team, 2020), both of which can calculate toxicity scores for a set of six toxic attributes. Scores range from 0 (least toxic) to 1 (most toxic). For each attribute, we calculate the proportion of texts whose toxicity scores exceed a 0.5 threshold.
4.1.3. Implementation Details
To evaluate our attack methodology, we employ the Vicuna-13B version of MiniGPT-4 (Zhu et al., 2023). It is built on the frozen Vicuna-13B (Chiang et al., 2023) backbone model, an aligned LLM derived from LLaMA (Touvron et al., 2023). For the experimental setup, we adopt the configuration utilized in VAJM (Qi et al., 2023), specifying a batch size $b$ of 8, a step size $\alpha$ of 1, and an iteration count $N_{1}$ of 5000. Furthermore, we set the image-text attack ratio $r$ to 10, the length of the adversarial text suffix to 20 tokens, the number of candidates $n$ to 250, and the iteration count $N_{2}$ to 2000 by default. We use a single A100 GPU with 80GB of memory for all experiments.
4.2. Comparison With Unimodal Attacks
@@
Table 1. Comparison of Train ASR (%) and Test ASR (%) on Advbench’s (Zou et al., 2023) harmful behaviors. GCG-V refers to the visual version of GCG.
@@
Table 2. Comparison of average ASR (%) measured on the VAJM (Qi et al., 2023) evaluation set across 4 categories of harmful instructions.
ASR(%) Identity Attack Disinfo Violence/Crime X-risk
Without Attack 30.8 53.3 57.3 33.3
GCG (Zou et al., 2023) 49.2 48.9 57.3 40.0
VAJM (Qi et al., 2023) 81.5 82.2 85.3 60.0
GCG-V 66.2 64.4 84.0 6.7
UMK (Ours) 87.7 95.6 98.7 46.7
Table 3. Percentages (%) of outputs that display specific toxic attributes evaluated by the Perspective API and Detoxify classifier (Hanu and Unitary team, 2020). These outputs are generated on the ‘challenging’ subset from RealToxicityPrompts benchmark (Gehman et al., 2020). ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
@@
Table 4. Comparison of Train ASR (%) and Test ASR (%) for ablation studies on Advbench’s (Zou et al., 2023) harmful behaviors.
@@
To further broaden our evaluation, we utilize the challenging subset of RealToxicityPrompts benchmark (Gehman et al., 2020), which includes 1,225 text prompts designed to trigger toxic continuations. Given that this is a text continuation task, we exclude the GCG method (Zou et al., 2023), which focuses on generating text suffixes, from our comparison. Similarly, for our approach, we utilize only the adversarial image prefix. We pair each adversarial image with text prompts from the dataset as input and then measure the toxicity of the outputs using both the Perspective API and the Detoxify classifier [31], each capable of calculating toxicity scores for a set of six toxic attributes. For each attribute, we calculate the ratio of the generated texts with scores exceeding the threshold of 0.5. As shown in Table 3, VAJM (Qi et al., 2023) and GCG-V display different performances across two evaluation methods. Under the Perspective API, GCG-V achieves a 71.98% toxicity rate, higher than VAJM’s 67.17%. Conversely, with the Detoxify classifier (Hanu and Unitary team, 2020), VAJM records a 61.96% rate, exceeding GCG-V’s 56.34%. Although VAJM achieves results superior to those of our method in certain categories such as Identity Attacks, this is attributable to its optimization objective being primarily focused on generating harmful statements against gender, race, and humanity. In contrast, our method’s optimization objectives are more comprehensive. Although our method only utilizes adversarial image prefixes in this experiment, it achieves overall best results, recording Any* toxicity rates of 76.98% and 68.70% under the Perspective API and Detoxify classifier tests, respectively. These results significantly outperform the previous best rates of 71.98% from GCG-V and 61.96% from VAJM. This demonstrates the effectiveness of our dual optimization objective strategy: even though the second phase is aimed at guiding the model to generate affirmative responses, it successfully enhances the toxic semantics of the adversarial image. Moreover, the joint attack of the text-image multimodalities may have uncovered a more optimal solution space for the adversarial image.
Figure 4. Comparative Analysis of Loss Before and After Text Attack. The X-axis represents ‘Steps’, while the Y-axis denotes ‘Loss’.
4.3. Ablation Studies
@@
4.4. Understanding the Role of Text Attack in Multimodal Strategy
@@
Figure 5. Overview of the image-image attack strategy. We adopt the same dual optimization objectives as used in the text-image attack.
@@
Figure 6. Examples of challenging attacks demonstrating our text-image multimodal strategy. The approach effectively bypasses MiniGPT-4’s (Zhu et al., 2023) safety mechanisms, generating harmful content including messages promoting self-harm or suicide and derogatory comparisons of politicians to Nazis.
In this subsection, we explore the role of text in the text-image multimodal attack strategy. Figure 4 illustrates the variance in batch loss before and after the text attack throughout the execution of our attack methodology. For ease of analysis, we employ linear regression models to fit the curve. Observations at different steps indicate a significant reduction in batch loss after the text attack. As the number of steps increases, an overall downward trend in loss can be observed, demonstrating the efficacy of our method. Notably, even in the later stages of the attack, the reduction in batch loss before and after the text attack remains evident, emphasizing the pivotal contribution of text in amplifying the effectiveness of our strategy.
@@
Table 5. Comparative analysis of average ASR (%) between image-image and text-image attacks on the VAJM (Qi et al., 2023) evaluation set.
ASR(%) Identity Attack Disinfo Violence/Crime X-risk
Image-Image Attack 78.5 68.9 85.3 6.7
Text-Image Attack 87.7 95.6 98.7 46.7
To better understand the role of text in our proposed text-image multimodal attack strategy, we introduce an image-image attack for comparative analysis. As illustrated in Figure 5, during the attack, we input two adversarial images and optimized them simultaneously. In the case of MiniGPT-4 (Zhu et al., 2023), the adversarial images we input each occupies 32 tokens, while the adversarial text suffix proposed in our attack utilizes only 20 tokens. Given that the textual space is discrete and denser compared to the visual space, adversarial attacks in the textual domain are generally more demanding. Intuitively, one might anticipate that image-image attacks, which are optimized over a larger volume of tokens, would yield superior results. However, this is not the case. In Table 5, we report the average ASR for both the image-image attack and the proposed text-image attack on the VAJM (Qi et al., 2023) evaluation set. It is evident that the efficacy of the image-image attack is significantly inferior to that of our attack. We observe that the image-image attacks are more likely to trigger the model to repeat affirmative responses twice or to cease output after the affirmative response. We believe this is because the optimization objectives for both adversarial images in the second phase are to maximize the probability of the model’s affirmative responses. In this context, the semantics of the two images tend to overshadow that of the text input. Consequently, during the testing phase, two adversarial images containing the same semantics are more likely to cause the model to duplicate affirmative reactions or only produce affirmative responses. On the other hand, this phenomenon can also be attributed to the training strategies employed by VLMs. Since VLMs are typically trained on image-text pairs, they tend to experience performance degradation when encountering out-of-distribution inputs such as image-image-text trios.
5. Conclusion
@@
Acknowledgements.
This work was supported by the National Key R&D Program of China (Grant No. 2022ZD0160103) and the National Natural Science Foundation of China (Grant No. 62276067).